OpenStudy (ilovebmth1234):

If you are given the lengths of two sides of a right triangle and do not know the angle measures of any angle besides the right angle, the way to find the third side length of the triangle is by using the Pythagorean theorem. True False

OpenStudy (cursedwizard):

that is correct because \(a^2+b^2=c^2\)

OpenStudy (nino_jim):

true a2+b2=c2

OpenStudy (vhtran):

true

OpenStudy (ilovebmth1234):

thanks guys @CursedWizard could you give @Nino_jim a medal too for helping too? i like to play it fair on here

OpenStudy (vhtran):

did it =^.^=

OpenStudy (cursedwizard):

oke :)